# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ XII.1 Definition and Implementation
A new formulation of perturbation expansion [110,111], which we designate as “Mean-Field-Perturbation Theory (MFPT)”, is naturally suggested in view of the above considerations by defining the perturbation about the AH as:
@@
An immediate consequence of this prescription , eqn.(157) is the following result:
@@
for all ``$n"$ and ``$g"$, by virtue of eqns.(154) and (156). This eqn. (158) ensures that the perturbation correction remains sub-dominant, in the sense of quantum average , for arbitrary values of `$g^{\prime}$ and `$n^{\prime}$ as demanded (see condition-(b) mentioned above, i.e.
@@
(Here,the quantum average of an operator A is defined as: $\langle A\rangle\equiv\langle\phi_{n}|A|\phi_{n}\rangle)$.
Furthermore, eqn. (158) has a direct consequence that the first-order perturbation correction in the Rayleigh-Schrö dinger-perturbation series (RSPS) vanishes identically for all `$n^{\prime}$ and `$g^{\prime}$:
@@
It is shown later that these two general outcomes of eqn.(158) have important bearing on the properties of the perturbation-series (PS) in MFPT, which is defined analogous to eqn.(1) as follows:
@@
(hence-forth, we do not display the $n$-dependence of the various quantities, for notational convenience)
For the study of convergence properties of the above series, one needs to compute the energy-corrections $E_{k}(g)$ to arbitrary order `$k^{\prime}$. However, since the above series is not a power- series-expansion in $g$, we resort to the well-known [109] recipe of introducing an auxiliary, dummy parameter denoted as $\eta$, for generating a power series in this parameter (chosen real) and project out the `$k^{\prime}$-th order correction, $E_{k}(g)$ by the following procedure: consider an associated-Hamiltonian (ASH), $\bar{H}$ given by
@@
${}~{}~{}~{}~{}~{}~{}~{}\bar{E}(\eta,g)=\sum_{k=0}^{\infty}\eta^{k}E_{k}(g)$ (166)
The `$k^{\prime}$-th perturbation correction $E_{k}(g)$ being the same as appearing in eqn.(161), can then be identified from the above equation as the coefficient of $\eta^{k}$ before setting the limit, $\eta\rightarrow 1$.
It may be emphasized that the above procedure is merely an intermediate book-keeping device to project out the $E_{k}(g)$ appearing in eqn. (161). Apart from achieving this goal, the above formalism plays no other role here. In particular, the final results are independent of the dummy-variable $\eta$ since, by construction, $H(g)=\bar{H}(\eta=1,g)$, $E(g)=\bar{E}(\eta=1,g)$, etc.
We determine the corrections $E_{k}(g)$ to arbitrary order `$k^{\prime}$ in MFPT, using the recursion relations derived from the application of the “hyper-virial theorem (HVT)” and the “Feynman-Hellman theorem (FHT)” to the Hamiltonian given by eqn.(163). This method is distinguished from other methods in generating exact values for perturbation corrections to an arbitrary order whereas other methods are afflicted by inevitable errors such as neglecting sub-asymptotic corrections, truncation-errors etc.(Moreover, the ease of implementation of recursive evaluation through widely available symbolic-computation, is an additional practical advantage of this method.) Here, we follow procedure analogous to that in ref.[9] used in the context of SFPT for the QAHO, with due generalization for extending the application to MFPT as outlined below in the next section.
@@ XIII.1 Tools: The Feynman-Hellman Theorem (FHT) and the Hyper-Virial Theorem (HVT)
In many cases of physical interest, the simultaneous application of the Feynman-Hellman Theorem and the hyper-virial theorem, renders the computation of perturbation corrections $E_{k}$ much easier without the necessity of summing infinite (sub)-series at each order. Furthermore, each term: $E_{k}$ can be precisely computed (with zero error) as inevitably occurs due to finite truncation of the conventional perturbation series at each order. Moreover, precise computation of $E_{k}$ to an arbitrary high-value of ‘$k$’ can be exactly computed through (system-specific-) recursion relations. For the AHI, the program was carried out by Swenson and Danforth [81] and elaborated in texts e.g. by Fernandez[9]. We discuss these theorems in turn, as follows.
@@ XIII.1.1 The Hyper-virial Theorem(HVT)
where,$\hat{H}$ is the (hermitian) Hamiltonian of the system and $[\hat{A},\hat{B}]$ stands for the commutator of operators, $\hat{A}$ and $\hat{B}$ and $\hat{W}\equiv\hat{W}{(\hat{x},\hat{p})}$ is any arbitrary, time independent operator function of $\hat{x}$ and $\hat{p}$ in the relevant Hilbert space. Here, we work in the co-ordinate space- representation such that $\hat{p}~{}=~{}-i\frac{\partial}{\partial{x}}$ and $\hat{x}~{}=~{}x$ are the momentum and position operators.
The above eqn.(165) is a consequence of the Energy-eigenvalue equation(and its hermitian conjugate) as appropriate for the time-independent ( ’Stationary’) situation and can be verified directly using :
@@ XIII.1.2 The Hellmann-Feynman Theorem(HFT)
Consider when the Hamiltonian, ${H}$ of the system depends on a parameter $\lambda$ , apart from the dynamic-variables of the system under consideration. (This situation is usually the case since $\lambda$ could be the system- charge, mass, a coupling-strength of interaction or even a dummy parameter introduced to define perturbation theory etc.) Then it follows that the eigenvalues and eigenfunctions of ${H}$ will also depend on that parameter.
@@
The partial derivative of the total energy with respect to an parameter occurring in the Hamiltonian of the system is equal to the expectation value of the partial derivative of the Hamiltonian with respect to the same parameter, as given by the following equation:
The proof (see,e.g.[9]) of the theorem is a straight forward application of the eigenvalue equation and the hermitian-property of the Hamiltonian for a stationary system , as given by:
@@
( In the above context, it may be relevant to remark that: (i) there are generalizations of the (FHT) to its ‘off-diagonal’ version, see ref.[106] and that (ii) the off-diagonal-FHT finds important applications e.g. in reproducing the RS-Perturbation theory without the use of the unperturbed wave-function).
@@ XIII.2 Tools : The Method of Swenson and Danforth
The method of Swenson and Danforth [ 81], utilizes the hyper-virial and the Hellmann-Feynman theorems to derive the perturbation-correction for the anharmonic oscillators in the SFT-framework. A lucid description of the same is provided in ref.[9] , which we closely follow in this subsection and present an overview of the method.
To start with, we consider the Hamiltonian for a non-relativistic, stationary system in one space-dimension given by:   $\hat{H}=\hat{p}^{2}{/}{2m}+V(\hat{x})$, and use the co-ordinate representation such that $\hat{x}=x$ and $\hat{p}=-i\hbar\partial{/}\partial x$.
In the next step, we explore the consequence of the HVT by choosing a suitable linear operator, $\hat{W}$ depending up on $\hat{x}$ and $\hat{p}$ Following ref. [9] , the following choice is made :
@@
where, $f(x)$ and $g(x)$ are arbitrary, ’regular’ (differentiable) functions to start with but allowing to be chosen conveniently for the system under study. Considering the one-dimensional Hamiltonian operator as given above, i.e.   $\hat{H}=\hat{p}^{2}{/}{2m}+V(\hat{x})$,   and using eqn.(169) in eqn.(165), one obtains:
@@
Using the basic commutation relation,   $[\hat{x},\hat{p}]=-i\hbar$   and the result: $[\hat{p},f(x)]=i\hbar f^{\prime}$ (where denotes derivative with respect to ${}x~{}$) and employing the standard manipulation techniques for $commutators$, one obtains the following result for $A$:
@@
Similarly, the $B$ term can be evaluated as
@@
$02E\langle{f^{\prime}}\rangle-2\langle{f^{\prime}}V\rangle-\langle f{V^{\prime}}\rangle+\left(\frac{\hbar^{2}}{4m}\right)\langle{f^{\prime\prime\prime}}\rangle~{}=~{}$ 0 . (181)
In the above equation, the notation is as follows: $E$ = energy eigenvalue; $f=f(x)$ is an arbitrary differentiable function which can be conveniently chosen for the particular problem at hand,“prime(s)” denote differentiation and $\langle A\rangle\equiv\langle\psi|A|\psi\rangle$ for an operator A, with $|\psi\rangle$ denoting the (normalized) eigenfunction of $H$.
It may be noted here that: (a) alternative derivation of equation (179) can be achieved, e.g. [9] by using the Jacobi identity for commutators. (b) This equation along with use of the Hellmann-Feynman Theorem, will be shown (later in this Section) to lead to recursive derivation of exact perturbation-corrections at ‘arbitrary’ order. (c) The yet-arbitrary function ‘$f$’ can be chosen according to the specific functional form for the potential-function: $V(x)$.
In the following subsection, we illustrate the SD-Method by applying it to the case of the general an-harmonic interaction in one-space dimension.
@@ XIII.3 The Swenson-Danforth Method (SDM) for the An-Harmonic Interactions (AHI)
$02E\langle{f^{\prime}}\rangle-2\langle{f^{\prime}}V\rangle-\langle f{V^{\prime}}\rangle+(\frac{1}{4})\langle{f^{\prime\prime\prime}}\rangle~{}=~{}$ 0 . (184)
@@
where’$j$’= 0,1,2,3,4….. Computing ${f^{\prime}}$, ${V^{\prime}}$, ${f^{\prime\prime\prime}}$, etc, substituting into eqn.(182) and simplifying , one obtains:
@@
where , $n=0,1,2,3,......$. (Henceforth, we suppress the label ${}^{\prime}n^{\prime}$ in $E$ and ${\cal E}$ for notational convenience.) Denote the eigenstates of $H_{SHO}$ by
@@
Recursive solution for $Y(j)$; $j$=0,1,2,…. can be obtained from eqn.(192)in terms of the known eigenvalues ${\cal E}$ Noting that: $Y(0)~{}=~{}1$ and $Y(j)\equiv 0$ for $j<0$ (by choice), we get
@@
This eqn.(198) thus relates the perturbation correction at order-$p$ in terms of the $X(K,p-1)$ , which can be evaluated recursively by the recurrence relation given by eqn.(187),again by expanding the relevant observables in power-series in $\lambda$ as given by eqn.(194) and reproduced below:
[…]
@@
+$+\left\{\frac{j(4j^{2}-1)}{4(j+1)}\right\}\sum_{i=0}^{\infty}\lambda^{i}X(j-1,i)$
@@
For, non-zero values of $\lambda$ i.e. for $i$=1,2,3,…. one gets from (eqn.202), the following recursion relation for $X(j+1,i)$ by the following procedure: Rewrite (202) by changing $i\rightarrow(i-1)$ in the last-term) leads to
@@
+$+\left\{\frac{(j-1)(4(j-1)^{2}-1)}{4j}\right\}X(j-2,i)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
@@
$E_{p}~{}=~{}\left(\frac{1}{p}\right)X(K,p-1),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{(198)}$ (209)
one can completely solve the above recursion relations for $E_{p}$; $p$=1,2,3,….; provided, the following boundary conditions are imposed,
@@
In respect of the recursion relation, eqn.206 the following observations are in order:
(a) it may be noted that the recursion terminates only when second argument of the last-term in eqn.(206) becomes zero due to the boundary conditions, eqn.(207). This implies that for evaluation of $E_{p}$, one needs to compute all coefficients $X(j,i)$ for $i=1,2,3,....(p-1)$; and for each fixed $i~{},~{}~{}j=1,2,3,......{(p-i)(K-1)+1}$ such that the “maximum” value for ‘${}j~{}$’ goes with “minimum” value of ‘${}i~{}$’. This specifies the set of values of ${}i,j~{}$ over which the recursion in eqn.(206) is to be carried out.
(b) The array of $X(j,i)$ required to solve the problem for computation of $E_{p}$ (i.e.$p$-th perturbation correction) can be generated by both numerical or symbolic implementation of computation. However, it soon becomes unwieldy in ’numerical’ implementation when $p\gg 1$ , due to large rounding-off errors and memory-restrictions. In ref.[9],symbolic implementation through “MAPLE” has been advocated, which we have followed and checked its suitability for $p\leq 200$, which suffices for most theoretical purposes (such as studies of large order behavior of SFPT).
(c) Sample results in SFPT for the QAHO were obtained by the recursion method, eqn.(206), which correctly reproduces [9] the results from WKB-asymptolic computation of Bender-Wu, Banks,Bender and Wu [88,100] for establishing the large-order-behavior confirming the Asymptotic nature and the predicted divergence in SFPT.
(d) More importantly, the“Hypervirial-Feynman-Hellman-Theorem-(HVT-FHT)” - Method, followed here, provides exact (i.e. no truncation) results at each order of perturbation theory.
XIV Application of MFPT to anharmonic-interactions
In the previous section, the method based upon the Hypervirial- and the Feynman-Hellman Theorems (HVT-FHT) for computing perturbation corrections in the “standard formulation of perturbation theory”(SFPT) was highlighted and its supremacy over the conventional method was demonstrated. The ease and simplicity of implementation of the same was established by reproduction of the results for perturbation corrections of arbitrary order in SFPT for the different cases of anharmonic oscillator obtained by other approaches.
In the following sub-sections, we carry out the application of the HVT-FHT method in the framework of the “Mean-Field Approximation Scheme” and the “Mean Field Perturbation Theory” (MFPT) to the case of the QAHO, SAHO and QDWO, which is among the main objectives of this work , in view of the inadequacies of SFPT stated earlier ( see, Sections -II, IV and XI).

### v3

@@ XII.1 Definition and Implementation
A novel formulation of perturbation expansion [110,111], termed as ”Mean-Field-Perturbation Theory (MFPT),” is naturally suggested in light of the aforementioned considerations by defining the perturbation around the AH as:
@@
An immediate consequence of this prescription, eqn. (157), is the following result:
@@
for all n``n" and g``g", by virtue of eqns. (154) and (156). This eqn. (158) ensures that the perturbation correction remains sub-dominant, in the sense of quantum average, for arbitrary values of $g^{\prime}$ and $n^{\prime}$ as required (see condition-(b) mentioned above, i.e.
@@
(Here, the quantum average of an operator A is defined as: $\langle A\rangle\equiv\langle\phi_{n}|A|\phi_{n}\rangle$). Furthermore, eqn. (158) has a direct consequence that the first-order perturbation correction in the Rayleigh-Schrödinger perturbation series (RSPS) vanishes identically for all $n^{\prime}$ and $g^{\prime}$:
@@
It is shown later that these two general outcomes of eqn. (158) have significant implications for the properties of the perturbation series (PS) in MFPT, which is defined analogously to eqn. (1) as follows:
@@
(Henceforth, we do not display the $n$-dependence of the various quantities, for notational convenience.)
For the study of convergence properties of the above series, one needs to compute the energy corrections $E_{k}(g)$ to arbitrary order $k^{\prime}$. However, since the above series is not a power series expansion in $g$, we resort to the well-known [109] recipe of introducing an auxiliary, dummy parameter denoted as $\eta$, for generating a power series in this parameter (chosen real) and project out the $k^{\prime}$-th order correction, $E_{k}(g)$ by the following procedure: consider an associated Hamiltonian (ASH), $\bar{H}$ given by
@@
$\bar{E}(\eta,g)=\sum_{k=0}^{\infty}\eta^{k}E_{k}(g)$ (166)
The $k^{\prime}-$th perturbation correction $E_{k}(g)$, as referenced in equation (161), can be discerned from the preceding equation as the coefficient of $\eta^{k}$ prior to setting the limit $\eta\rightarrow 1$. It is important to note that this procedure serves merely as an intermediate book-keeping mechanism to isolate the $E_{k}(g)$ appearing in equation (161). Beyond this purpose, the formalism does not serve any additional function in this context. Specifically, the final results are independent of the dummy variable $\eta$ since, by design, $H(g)=\bar{H}(\eta=1,g)$, $E(g)=\bar{E}(\eta=1,g)$, and so forth. We compute the corrections $E_{k}(g)$ to an arbitrary order $k^{\prime}$ in MFPT, utilizing the recursion relations derived from the application of the ”hyper-virial theorem (HVT)” and the ”Feynman-Hellman theorem (FHT)” to the Hamiltonian specified in equation (163). This method is distinct from other approaches in that it generates exact values for perturbation corrections to any order, whereas other methods are prone to inevitable errors such as neglecting sub-asymptotic corrections and truncation errors. Additionally, the ease of implementing recursive evaluation through widely available symbolic computation is a practical advantage of this method. Here, we follow a procedure analogous to that in reference [9] used in the context of SFPT for the QAHO, with appropriate generalization for extending the application to MFPT as outlined in the subsequent section.
@@ XIII.1 Tools: The Feynman-Hellman Theorem (FHT) and the Hyper-Virial Theorem (HVT)
In many cases of physical interest, the concurrent application of the Feynman-Hellman Theorem and the hyper-virial theorem facilitates the computation of perturbation corrections $E_{k}$ without the necessity of summing infinite (sub)-series at each order. Furthermore, each term $E_{k}$ can be computed precisely (with zero error), as is inevitably the case due to finite truncation of the conventional perturbation series at each order. Moreover, precise computation of $E_{k}$ to an arbitrarily high value of ‘$k$’ can be achieved exactly through system-specific recursion relations. For the AHI, the program was executed by Swenson and Danforth [81] and elaborated in texts such as by Fernandez [9]. We discuss these theorems in turn, as follows.
@@ XIII.1.1 The Hyper-virial Theorem(HVT)
[…]
Then the above eqn.(165) follows as a consequence of the appropriate energy-eigenvalue equation(and its hermitian conjugate) for the time-independent ( ’Stationary’) situation, which can be verified directly using :
@@ XIII.1.2 The Hellmann-Feynman Theorem(HFT)
Consider a scenario where the Hamiltonian, H{H}, of a system is dependent on a parameter $\lambda$, in addition to the dynamic variables of the system under consideration. This situation is quite common, as $\lambda$ could represent the system’s charge, mass, a coupling strength of interaction, or even a dummy parameter introduced to define perturbation theory, among others. Consequently, the eigenvalues and eigenfunctions of H{H} will also be dependent on this parameter. The Hellmann-Feynman Theorem (HFT) is applicable in this context and yields significant results in perturbation theories without necessitating the use of wave functions, as discussed in ref. [9]. The theorem posits that:
The partial derivative of the total energy with respect to a parameter present in the Hamiltonian of the system is equivalent to the expectation value of the partial derivative of the Hamiltonian with respect to the same parameter, as expressed by the following equation:
@@
In this context, it is pertinent to note that: (i) there are generalizations of the (FHT) to its ’off-diagonal’ version, as referenced in [106], and (ii) the off-diagonal FHT has significant applications, such as in reproducing the RS-Perturbation theory without the use of the unperturbed wave function.
@@ XIII.2 Tools : The Method of Swenson and Danforth
The method of Swenson and Danforth [81] employs the hyper-virial and the Hellmann-Feynman theorems to derive the perturbation correction for anharmonic oscillators within the SFT framework. A comprehensive description of this method is provided in ref. [9], which we closely follow in this subsection to present an overview of the method.
[…]
In the next step, we explore the consequence of the HVT by choosing a suitable linear operator, $\hat{W}$ depending up on $\hat{x}$ and $\hat{p}$ . Following ref. [9] , the following choice is made :
@@
[…]
@@
Using the basic commutation relation,   $\hat{x},\hat{p}]=-i\hbar$   and the result: $\hat{p},f(x)]=i\hbar f^{\prime}$ (where denotes derivative with respect to $x~$) and employing the standard manipulation techniques for commutatorscommutators, one obtains the following result for $A$:
@@
Similarly, the $B$ - term can be evaluated as
@@
$2E\langle{f^{\prime}}\rangle-2\langle{f^{\prime}}V\rangle-\langle f{V^{\prime}}\rangle+\left(\frac{\hbar^{2}}{4m}\right)\langle{f^{\prime\prime\prime}}\rangle~=~0.$ (181)
In the above equation, the notation is as follows: $E$ represents the energy eigenvalue; $f=f(x)$ is an arbitrary differentiable function that can be conveniently chosen for the particular problem at hand, ”prime(s)” denote differentiation, and $\langle A\rangle\equiv\langle\psi|A|\psi\rangle$ for an operator $A$, with $|\psi\rangle$ denoting the (normalized) eigenfunction of $H$.
It is noteworthy that: (a) an alternative derivation of equation (179) can be achieved, e.g., [9] by using the Jacobi identity for commutators. (b) This equation, along with the use of the Hellmann-Feynman Theorem, will be shown (later in this Section) to lead to a recursive derivation of exact perturbation corrections at arbitrary order. (c) The yet-arbitrary function $f$ can be chosen according to the specific functional form for the potential function: $V(x)$. In the following subsection, we illustrate the SD-Method by applying it to the case of the general anharmonic interaction in one spatial dimension.
(Up to this point on 25 Aug 2025)
@@ XIII.3 The Swenson-Danforth Method (SDM) for the An-Harmonic Interactions (AHI)
$2E\langle{f^{\prime}}\rangle-2\langle{f^{\prime}}V\rangle-\langle f{V^{\prime}}\rangle+(\frac{1}{4})\langle{f^{\prime\prime\prime}}\rangle~=~0.$ (184)
@@
where’jj’= 0,1,2,3,4….. Computing $f^{\prime}$}, $V^{\prime}$}, $f^{\prime\prime\prime}$}, etc, substituting into eqn.(182) and simplifying , one obtains:
@@
[…]
@@
[…]
@@
This eqn.(198) thus relates the perturbation correction at order-$p$ in terms of the X(K,p1)X(K,p-1) , which can be evaluated recursively by the recurrence relation given by eqn.(187),again by expanding the relevant observables in power-series in $\lambda$ as given by eqn.(194) and reproduced below:
$E\equiv E(\lambda)=\sum_{p=0}^{\infty}E_{p}\lambda^{p},X(j,\lambda)\equiv\sum_{i=0}^{\infty}\lambda^{i}X(j,i).~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\mbox{(194)}$
@@
$+\left\{\frac{j(4j^{2}-1)}{4(j+1)}\right\}\sum_{i=0}^{\infty}\lambda^{i}X(j-1,i)$
@@
For, non-zero values of $\lambda$ i.e. for ii=1,2,3,…. one gets from (eqn.202), the following recursion relation for $X(j+1,i)$ by the following procedure: Rewrite (202) by changing $i\rightarrow(i-1)$ in the last-term) leads to
@@
$+\left\{\frac{(j-1)(4(j-1)^{2}-1)}{4j}\right\}X(j-2,i)$~~~~~~~~~~~~
@@
$E_{p}~=~\left(\frac{1}{p}\right)X(K,p-1),~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~{(198)}$ (209)
one can completely solve the above recursion relations for $E_{p}$; pp=1,2,3,….; provided, the following boundary conditions are imposed,
@@
Regarding the recursion relation, equation (206), the following observations are pertinent: (a) The recursion terminates only when the second argument of the last term in equation (206) becomes zero due to the boundary conditions, equation (207). This implies that for the evaluation of $E_{p}$, one needs to compute all coefficients $X(j,i)$ for $i=1,2,3,\ldots,(p-1)$; and for each fixed $i$, $j=1,2,3,\ldots,(p-i)(K-1)+1$, such that the ”maximum” value for ‘$j$’ corresponds with the ”minimum” value of ‘$i$’. This specifies the set of values of $i,j$ over which the recursion in equation (206) is to be carried out. (b) The array of $X(j,i)$ required to solve the problem for the computation of $E_{p}$ (i.e., $p$-th perturbation correction) can be generated by both numerical or symbolic implementation of computation. However, it soon becomes unwieldy in ’numerical’ implementation when $p\gg 1$, due to large rounding-off errors and memory restrictions. In reference [9], symbolic implementation through ”MAPLE” has been advocated, which we have followed and verified its suitability for $p\leq 200$, which suffices for most theoretical purposes (such as studies of large order behavior of SFPT). (c) Sample results in SFPT for the QAHO were obtained by the recursion method, equation (206), which correctly reproduces [9] the results from asymptotic computation of Bender-Wu, Banks, Bender, and Wu [88,100] for establishing the large-order behavior, confirming the asymptotic nature and the predicted divergence in SFPT. (d) More importantly, the ”Hypervirial-Feynman-Hellman-Theorem-(HVT-FHT)” method, followed here, provides exact (i.e., no truncation) results at each order of perturbation theory.
XIV Application of MFPT to Anharmonic Interactions
In the preceding section, the method based on the Hypervirial and Feynman-Hellman Theorems (HVT-FHT) for calculating perturbation corrections within the ”standard formulation of perturbation theory” (SFPT) was emphasized, demonstrating its superiority over conventional methods. The ease and simplicity of its implementation were established by reproducing results for perturbation corrections of arbitrary order in SFPT for various cases of anharmonic oscillators, as obtained by other approaches. In the subsequent subsections, we apply the HVT-FHT method within the framework of the ”Mean-Field Approximation Scheme” and the ”Mean Field Perturbation Theory” (MFPT) to the cases of the Quartic Anharmonic Oscillator (QAHO), Sextic Anharmonic Oscillator (SAHO), and Quartic Double Well Oscillator (QDWO), which constitute the primary objectives of this study, given the inadequacies of SFPT previously discussed (see Sections II, IV, and XI).